# Least multivariate Chebyshev polynomials on diagonally-determined domains

### v2

@@ Abstract.
We consider a new multivariate generalization of the classical monic (univariate) Chebyshev polynomial that minimizes the uniform norm on the interval $[-1,1]$. Let $\Pi^{*}_{n}$ be the subset of polynomials of degree at most $n$ in $d$ variables, whose homogeneous part of degree $n$ has coefficients summing up to $1$. The problem is determining a polynomial in $\Pi^{*}_{n}$ with the smallest uniform norm on a domain $\Omega$, which we call a least Chebyshev polynomial (associated with $\Omega$). Our main result solves the problem for $\Omega$ belonging to a non-trivial class of domains that we call diagonally-determined, and establishes the remarkable result that a least Chebyshev polynomial can be given via the classical, univariate, Chebyshev polynomial. In particular, the solution can be independent of the dimension. Diagonally-determined domains include centered balls in $\mathbb{R}^{d}$ in any norm, but can be non-convex and highly irregular. We also introduce a computational procedure, based on semidefinite programming hierarchies, to detect if a given semi-algebraic set is diagonally-determined.
@@
Least polynomial, Chebyshev polynomial, best approximation, extremal signature
@@ 1. Introduction
Among its numerous properties, the Chebyshev polynomial $T_{n}(x)=\cos(n\arccos x)$ provides a solution for the best approximation to the monomial $x^{n}$ on the interval $[-1,1]$ in the uniform norm. More precisely, the polynomial
@@
of degree $n-1$ ($n\in\mathbb{N}$) is the best polynomial of approximation to $x^{n}$ on $[-1,1]$; that is
@@
with $\Pi_{n-1}$ being the vector space of univariate polynomials of degree at most $n-1$. In other words, the monic Chebyshev polynomial $x^{n}-q^{*}_{n}$ is the least polynomial in the sense that it has the least uniform norm among all monic polynomials of degree $n$.
There have been multiple extensions of Chebyshev polynomials to multivariate settings from different angles. From the point of view of approximation, an immediate generalization is finding the best approximation to monomials [1, 2, 5, 3, 9, 7, 11, 15, 16, 17]. Namely, for $d>1$, ${\alpha}\in{\mathbb{N}}^{d}$ and $|{\alpha}|\coloneqq\sum_{i=1}^{d}\alpha_{i}=n$, we consider the problem
@@
where $\Pi^{d}_{n-1}$ denotes the real polynomials of total degree at most $n-1$ in the variables ${\mathbf{x}}=(x_{1},\ldots,x_{d})$, and where we define the monomial ${\mathbf{x}}^{\alpha}\coloneqq x_{1}^{\alpha_{1}}\ldots x_{d}^{\alpha_{d}}$. Since we will deal with monomials of degree $n$ throughout the paper, we introduce the notation
@@
where ${\mathbb{N}}_{0}={\mathbb{N}}\cup\{0\}$ denotes the nonnegative integers. Thus ${\mathbb{N}}^{d}_{n}$ indexes all monomials of degree exactly $n$ in $d$ variables.
Problem (1.2) can be regarded as a natural multivariate generalization of (1.1), where $\Omega\subset{\mathbb{R}}^{d}$ is a subset of ${\mathbb{R}}^{d}$. While the interval $[-1,1]$ is a prototype of a compact connected set of the real line, there is no ‘prototype’ in higher dimensions for such a set. In the literature, this problem has been studied primarily on a few special regular domains. In two variables, the problem (1.2) is solved for the square, the disk, and the isosceles right triangle [5, 3, 7, 11, 15]. While the solution on the square can be extended to the cube for $d>2$, the problem is solved only for a few cases, mostly monomials of lower degrees, on the ball and the simplex [1, 2, 16]. Moreover, the existing examples indicate an increasing complexity, so much so that it does not appear possible to find an analytic solution even for these regular domains.
Recently, in [4] we have proposed to investigate (1.2) for various choices of $\Omega\subset{\mathbb{R}}^{d}$ and ${\alpha}\in{\mathbb{N}}^{d}_{n}$ by combining analytical tools with numerical tools from optimization (and notably the so-called moment-SOS hierarchy). During this study, we have encountered an optimization problem that has initiated a change of view: Namely, instead of studying the best polynomial of approximation to monomials, we can study the least polynomial instead. While the two concepts are identical in one variable, they can be quite different in higher dimensions, as seen from the definition below.
@@
Let $\Omega$ be a domain in ${\mathbb{R}}^{d}$. We consider the optimization problem
@@
If it exists, we call a minimizer $P^{*}\in\Pi^{*}_{n}$ of (1.3) a least Chebyshev polynomial of degree $n$ on the domain $\Omega$.
For $d=1$, there is only one monomial of degree $n$. In the case $d>1$, every element of $\Pi^{*}_{n}$ is ‘monic’ and the monomial ${\mathbf{x}}^{\alpha}$ in (1.1) is only one among many possible choices in $\Pi^{*}_{n}$. However, rather than approximating a fixed monomial by polynomials of lower degree, the problem (1.3) requires finding a polynomial that has the least norm among all polynomials in $\Pi_{n}^{*}$. As far as we are aware, this polynomial has not been considered in the literature.
@@
The main purpose of this paper is to report our findings on the optimization problem (1.3). It turns out, much to our surprise, that the problem (1.3) can be solved analytically for a fairly general family of domains $\Omega$ in ${\mathbb{R}}^{d}$ for all $d\geq 2$. This family of domains will be referred to as diagonally-determined.
@@
The paper is organized as follows. We start by defining and describing several examples of diagonally-determined domains in Section 2. Our main results are presented in Section 3, where we describe the least polynomial for a diagonally-determined domain. We discuss the dual problem of (1.3) in Section 4 and show that it too has a closed-form solution in the case of diagonally-determined sets. We also rephrase this result for the dual problem in the framework of extremal signatures. Finally, we consider the problem of deciding whether a given semi-algebraic set is diagonally determined in Section 5, and propose an algorithmic approach based on semidefinite programming hierarchies. We conclude the paper with a discussion of related problems and open questions in Section 6.
2. Diagonally-determined domains
We define the diagonal of a domain $\Omega\subset\mathbb{R}^{d}$ as the set ${\rm diag}(\Omega)\coloneqq\{t\in{\mathbb{R}}:t\mathbf{1}\in\Omega\}$, where $\mathbf{1}$ denotes the all-ones vector in $\mathbb{R}^{d}$. For two vectors ${\mathbf{x}},{\mathbf{y}}\in{\mathbb{R}}^{d}$, we use $\langle{\mathbf{x}},{\mathbf{y}}\rangle$ for their standard inner product.
@@ 2. Diagonally-determined domains
Importantly, Definition 2.1 covers domains that could be non-convex, non-compact, and even highly irregular. Figure 1 gives an example of a non-convex diagonally-determined set in ${\mathbb{R}}^{2}$.
$x_{1}$$x_{2}$001111222233334444111122223333
@@
We will now show that balls in ${\mathbb{R}}^{d}$ (in any norm), centered at the origin, are examples of diagonally-determined sets.
@@
[…]
where $|{\mathbf{x}}|=(|x_{1}|,\ldots,|x_{d}|)$, then the vector ${\mathbf{v}}$ may be assumed to be entrywise-nonnegative, denoted by  ${\mathbf{v}}\geq\mathbf{0}$.
@@
We prove the statement for unit balls; the required result then follows from Lemma 2.2.
[…]
Let ${\mathbf{v}}={\mathbf{u}}/\|\mathbf{1}\|$, where ${\mathbf{u}}\in{\mathbb{R}}^{d}$ is ‘dual’ to $\mathbf{1}$, in the sense that $\|{\mathbf{u}}\|_{*}=1$ and $\langle{\mathbf{u}},\mathbf{1}\rangle=\|\mathbf{1}\|$. We then have ${\rm diag}(\Omega)=[a,b]$ with $a=-1/\|\mathbf{1}\|$ and $b=1/\|\mathbf{1}\|$, $\langle{\mathbf{u}}/\|\mathbf{1}\|,\mathbf{1}\rangle=1$, and $|\langle{\mathbf{u}}/\|\mathbf{1}\|,{\mathbf{x}}\rangle|\leq(1/\|\mathbf{1}\|)\|{\mathbf{u}}\|_{*}\|{\mathbf{x}}\|\leq 1/\|\mathbf{1}\|=b$, so that $\langle{\mathbf{u}}/\|\mathbf{1}\|,{\mathbf{x}}\rangle\in[a,b]$ for all ${\mathbf{x}}\in\Omega$.
[…]
$\|{\mathbf{u}}\|_{*}\geq\||{\mathbf{u}}|\|_{*}\quad\forall{\mathbf{u}}\in\mathbb{R}^{d}.$
Thus, the vector ${\mathbf{u}}$ ‘dual’ to $\mathbf{1}$ in the above construction is component-wise nonnegative. Indeed, if not, then
$1=\|{\mathbf{u}}\|_{*}\geq\||{\mathbf{u}}|\|_{*}\geq\langle|{\mathbf{u}}|,\mathbf{1}/\|\mathbf{1}\|\rangle>\langle{\mathbf{u}},\mathbf{1}/\|\mathbf{1}\|\rangle=1,$
a contradiction. Subsequently, since ${\mathbf{v}}={\mathbf{u}}/\|\mathbf{1}\|$, we have ${\mathbf{v}}\geq\mathbf{0}$, as required. ∎
@@
[…]
@@
Finally, the diagonally-determined property is not invariant under translation in general. The exception is a translation by a constant multiple of the all-ones vector. We state this observation as a lemma for later use. The proof again follows immediately from Definition 2.1 and is omitted.
@@
Fix $\alpha\in\mathbb{R}$. Then a given $\Omega\subset\mathbb{R}^{d}$ is diagonally determined with $\mathrm{diag}(\Omega)=[a,b]$ and vector ${\mathbf{v}}$ if and only if its translation $\Omega+\alpha\mathbf{1}$ is diagonally determined with diagonal $[a+\alpha,b+\alpha]$ and vector ${\mathbf{v}}$.
3. Least Chebyshev polynomial for a diagonally-determined domain
In this section, and with $\Omega$ being a domain in ${\mathbb{R}}^{d}$, we consider the problem of finding a least polynomial $P^{*}\in\Pi_{n}^{*}$ such that
@@ 3. Least Chebyshev polynomial for a diagonally-determined domain
$P({\mathbf{x}})=\sum_{{\alpha}\in{\mathbb{N}}^{d}_{n}}a_{\alpha}\,{\mathbf{x}}^{\alpha}+Q({\mathbf{x}})\,\quad\text{with }\quad\sum_{{\alpha}\in{\mathbb{N}}^{d}_{n}}a_{\alpha}=1\,\text{and }\>Q\in\Pi^{d}_{n-1}.$
Then $P(t,\ldots,t)=t^{n}+q_{n-1}(t)$, where $q_{n-1}(t)=Q(t,\ldots,t)$ is a univariate polynomial of degree at most $n-1$ in the single variable $t$. Hence,
@@
where we have used a classical result in one variable, and equality in the last inequality is attained by choosing $q_{n-1}$ such that $t^{n}+q_{n-1}(t)$ is the rescaled Chebyshev polynomial in the right-hand side of (3.2).
@@
The minimal value in (3.1) does not depend explicitly on the dimension $d$, but the diagonal $[a,b]$ may. For example, if $\Omega$ is the unit Euclidean ball in ${\mathbb{R}}^{d}$, then $[a,b]=[-1/\sqrt{d},1/\sqrt{d}]$.
@@
[…]
@@
Returning to Example 2.5, where $\Omega$ is the unit ball for the owl norm, we have $[a,b]=[-1/W,1/W]$ and ${\mathbf{v}}={\mathbf{w}}/W$, so that, by Theorem 3.1, the least Chebyshev polynomial of degree $n$ becomes
@@
We point out that the requirement $\sum_{{\alpha}\in{\mathbb{N}}^{d}_{n}}a_{\alpha}=1$, is susceptible to an affine change of variables other than a translation. For instance, the triangle ^$\hat{\triangle}^{2}=\{\,{\mathbf{x}}:x_{1}\geq 0,x_{2}\geq 0,x_{1}+x_{2}\leq 1\,\}$, is a mirror image of $\triangle^{2}$ under $x_{2}\mapsto 1-x_{2}$. However, under the affine change of variables $x_{2}\mapsto 1-x_{2}$, the leading monomial $a_{0}x_{1}^{2}+a_{1}x_{1}x_{2}+a_{2}x_{2}^{2}$ of a polynomial in $\Pi_{2}^{*}$ becomes, $a_{0}x_{1}^{2}-a_{1}x_{1}x_{2}+a_{2}x_{2}^{2}$, which is no longer an element of $\Pi_{2}^{*}$.
@@
$\Omega_{1}$ ::=:= $\{(x,x)\;|\;x\in[-1,1]\}$
$\Omega_{2}$ ::=:= $\{(0,x)\;|\;x\in[-\sqrt{2},\sqrt{2}]\}$
$\Omega_{3}$ ::=:= $\{(x,-x)\;|\;x\in[-1,1]\}.$

### v3

# Least multivariate Chebyshev polynomials on diagonally-determined sets

@@ Abstract.
We consider a new multivariate generalization of the classical monic (univariate) Chebyshev polynomial that minimizes the uniform norm on the interval [1,1][-1,1]. Let $\Pi^{*}_{n}$ be the subset of polynomials of degree at most $n$ in $d$ variables, whose homogeneous part of degree $n$ has coefficients summing up to $1$. The problem is determining a polynomial in $\Pi^{*}_{n}$ with the smallest uniform norm on a set $\Omega\subset\mathbb{R}^{d}$, which we call a least Chebyshev polynomial (associated with $\Omega$). Our main result solves the problem for $\Omega$ belonging to a non-trivial class of sets that we call diagonally-determined, and establishes the remarkable result that a least Chebyshev polynomial can be given via the classical, univariate, Chebyshev polynomial. In particular, the solution can be independent of the dimension. Diagonally-determined sets include centered balls in $\mathbb{R}^{d}$ in any norm, but can be non-convex and even non-simply connected. We also introduce a computational procedure, based on semidefinite programming hierarchies, to detect if a given semi-algebraic set is diagonally-determined.
@@
Chebyshev polynomial, Chebyshev approximation, extremal signature, semidefinite programming, Lasserre hierarchy
@@ 1. Introduction
Among its numerous properties, the Chebyshev polynomial $T_{n}(x)=\cos(n\arccos x)$ provides a solution for the best approximation to the monomial $x^{n}$ on the interval [1,1][-1,1] in the uniform norm. More precisely, the polynomial
@@
of degree n1n-1 ($n\in\mathbb{N}$) is the best polynomial of approximation to $x^{n}$ on [1,1][-1,1]; that is
@@
with $\Pi_{n-1}$ being the vector space of univariate polynomials of degree at most n1n-1. In other words, the monic Chebyshev polynomial $x^{n}-q^{*}_{n}$ is the least polynomial in the sense that it has the least uniform norm among all monic polynomials of degree $n$.
There have been multiple extensions of Chebyshev polynomials to multivariate settings from different angles. From the point of view of approximation, an immediate generalization is finding the best approximation to monomials [1, 2, 5, 3, 9, 7, 11, 15, 16, 17]. To this end we replace the interval [1,1][-1,1] by a general set $\Omega\subset\mathbb{R}^{d}$, for $d>1$. Letting ${\alpha}\in{\mathbb{N}}^{d}$ and $|{\alpha}|\coloneqq\sum_{i=1}^{d}\alpha_{i}=n$, we consider the problem
@@
where $\Pi^{d}_{n-1}$ denotes the space of real polynomials of total degree at most n1n-1 in the variables ${\mathbf{x}}=(x_{1},\ldots,x_{d})$, and where we define the monomial ${\mathbf{x}}^{\alpha}\coloneqq x_{1}^{\alpha_{1}}\ldots x_{d}^{\alpha_{d}}$. Since we will deal with monomials of degree $n$ throughout the paper, we introduce the notation
@@
where ${\mathbb{N}}_{0}={\mathbb{N}}\cup\{0\}$ denotes the set of nonnegative integers. Thus ${\mathbb{N}}^{d}_{n}$ indexes all monomials of degree exactly $n$ in $d$ variables.
Problem (1.2) can be regarded as a natural multivariate generalization of (1.1), where $\Omega\subset{\mathbb{R}}^{d}$ is a subset of ${\mathbb{R}}^{d}$. While the interval [1,1][-1,1] is a prototype of a compact connected set of the real line, there is no ‘prototype’ in higher dimensions for such a set. In the literature, this problem has been studied primarily on a few special regular domains. In two variables, the problem (1.2) is solved for the square, the disk, and the isosceles right triangle [5, 3, 7, 11, 15]. While the solution on the square can be extended to the cube for $d>2$, the problem is solved only for a few cases, mostly monomials of lower degrees, on the ball and the simplex [1, 2, 16]. Moreover, the existing examples indicate an increasing complexity, so much so that it does not appear possible to find an analytic solution even for these regular domains.
Recently, in [4] we have proposed to investigate (1.2) for various choices of $\Omega\subset{\mathbb{R}}^{d}$ and ${\alpha}\in{\mathbb{N}}^{d}_{n}$ by combining analytical tools with numerical tools from optimization (and notably the so-called moment-SOS hierarchy [6]). During this study, we have encountered an optimization problem that has initiated a change of view: Namely, instead of studying the best polynomial of approximation to monomials, we can study the least polynomial instead. While the two concepts are identical in one variable, they can be quite different in higher dimensions, as seen from the definition below.
We emphasize that – even though our paper deals with the characterization of least polynomials on special sets – it is intrinsically connected to constructive approximation and numerical optimization due to the close links with the work in [4] and [6] respectively.
@@
Let $\Omega$ be a set in ${\mathbb{R}}^{d}$. We consider the optimization problem
@@
If it exists, we call a minimizer $P^{*}\in\Pi^{*}_{n}$ of (1.3) a least Chebyshev polynomial of degree $n$ on the set $\Omega$.
For $d=1$, there is only one monomial of degree $n$. In the case $d>1$, every element of $\Pi^{*}_{n}$ is ‘monic’ and the monomial ${\mathbf{x}}^{\alpha}$ in (1.2) is only one among many possible choices in $\Pi^{*}_{n}$. However, rather than approximating a fixed monomial by polynomials of lower degree, the problem (1.3) requires finding a polynomial that has the least norm among all polynomials in $\Pi_{n}^{*}$. As far as we are aware, this problem has not been considered in the literature.
@@
The main purpose of this paper is to report our findings on the optimization problem (1.3). It turns out, much to our surprise, that the problem (1.3) can be solved analytically for a fairly general family of sets $\Omega$ in ${\mathbb{R}}^{d}$ for all $d\geq 2$. This family of sets will be referred to as diagonally-determined.
@@
The paper is organized as follows. We start by defining and describing several examples of diagonally-determined sets in Section 2. Our main results are presented in Section 3, where we describe the least polynomial for a diagonally-determined set; see Theorem 3.1. We discuss the dual problem of (1.3) in Section 4 and show that it too has a closed-form solution in the case of diagonally-determined sets; see Theorem 4.8. We also rephrase this result for the dual problem in the framework of extremal signatures. Finally, we consider the problem of deciding whether a given semi-algebraic set is diagonally determined in Section 5, and propose an algorithmic approach based on semidefinite programming hierarchies. We conclude the paper with a discussion of related problems and open questions in Section 6.
2. Diagonally-determined sets
We define the diagonal of a set $\Omega\subset\mathbb{R}^{d}$ as the set ${\rm diag}(\Omega)\coloneqq\{t\in{\mathbb{R}}:t\mathbf{1}\in\Omega\}$, where $\mathbf{1}$ denotes the all-ones vector in $\mathbb{R}^{d}$. For two vectors ${\mathbf{x}},{\mathbf{y}}\in{\mathbb{R}}^{d}$, we use $\langle{\mathbf{x}},{\mathbf{y}}\rangle$ for their standard inner product.
@@ 2. Diagonally-determined sets
Importantly, Definition 2.1 covers sets that could be non-convex, non-compact, and not simply connected. Figure 1 gives an example of a non-convex diagonally-determined set in ${\mathbb{R}}^{2}$.
$x_{1}$$x_{2}$011223344112233
@@
We will now show that balls in ${\mathbb{R}}^{d}$ (in any norm), centered at the origin, are examples of diagonally-determined sets. To this end, we first need to recall some definitions.
Definition 2.3.
For a given norm $\|\cdot\|$ on $\mathbb{R}^{d}$, we use $\|\cdot\|_{*}$ to denote its dual norm, namely
We will also consider the special case of monotone norms.
Definition 2.4.
We call a norm $\|\cdot\|$ on $\mathbb{R}^{d}$ monotone, if
where $|{\mathbf{x}}|=(|x_{1}|,\ldots,|x_{d}|)$.
We may now show that any centered ball is diagonally-determined.
@@
where $\|\cdot\|$ denotes any norm on ${\mathbb{R}}^{d}$. Then $\Omega$ is a diagonally-determined set. If, in addition, the norm $\|\cdot\|$ is monotone,then the vector $\mathbf{v}$} may be assumed to be entrywise-nonnegative, denoted by  ${\mathbf{v}}\geq\mathbf{0}$.
@@
We prove the statement for unit balls; the required result then follows from Lemma 2.2. For a centered unit ball, one has ${\rm diag}(\Omega)=[a,b]$ with $a=-1/\|\mathbf{1}\|$ and $b=1/\|\mathbf{1}\|$.
Now let ${\mathbf{v}}={\mathbf{u}}/\|\mathbf{1}\|$, where ${\mathbf{u}}\in{\mathbb{R}}^{d}$ is ‘dual’ to $\mathbf{1}$, in the sense that $\|{\mathbf{u}}\|_{*}=1$ and $\langle{\mathbf{u}},\mathbf{1}\rangle=\|\mathbf{1}\|$. We then have $\langle{\mathbf{v}},\mathbf{1}\rangle=\langle{\mathbf{u}}/\|\mathbf{1}\|,\mathbf{1}\rangle=1$, as required. Moreover we have, for any ${\mathbf{x}}\in\Omega$,
$|\langle{\mathbf{v}},{\mathbf{x}}\rangle|=|\langle{\mathbf{u}}/\|\mathbf{1}\|,{\mathbf{x}}\rangle|\leq(1/\|\mathbf{1}\|)\|{\mathbf{u}}\|_{*}\|{\mathbf{x}}\|\leq 1/\|\mathbf{1}\|.$
Thus we see that $\langle{\mathbf{v}},{\mathbf{x}}\rangle\in[a,b]$ for all ${\mathbf{x}}\in\Omega$, proving that $\Omega$ is diagonally-determined.
Assume now that the norm $\|\cdot\|$ is monotone. We have
$\langle{\mathbf{u}},\mathbf{1}\rangle=\|\mathbf{1}\|\geq\|\mbox{sgn}({\mathbf{u}})\|\geq\langle{\mathbf{u}},\mbox{sgn}({\mathbf{u}})\rangle,$
where, for $i\in\{1,\ldots,d\}$,
$\mbox{sgn}({\mathbf{u}})_{i}=\left\{\begin{array}[]{rl}1&\mbox{if $u_{i}\geq 0$}\\ -1&\mbox{else.}\end{array}\right$.
This means that $\sum_{i=1}^{d}u_{i}\geq\sum_{i=1}^{d}|u_{i}|$, which holds if and only of ${\mathbf{u}}\geq\mathbf{0}$. Subsequently, since ${\mathbf{v}}={\mathbf{u}}/\|\mathbf{1}\|$, we have ${\mathbf{v}}\geq\mathbf{0}$.
@@
Assume $\|\cdot\|$ is a norm on ${\mathbb{R}}^{d}$ with dual norm $\|\cdot\|_{*}$, and that the following holds: there is an entrywise-nonnegative vector ${\mathbf{u}}\geq\mathbf{0}$ such that $\|{\mathbf{u}}\|_{*}=1$ and $\langle{\mathbf{u}},\mathbf{1}\rangle=\|\mathbf{1}\|$.
@@
Finally, the property of being diagonally-determined is not invariant under translation in general. The exception is a translation by a constant multiple of the all-ones vector. We state this observation as a lemma for later use. The proof again follows immediately from Definition 2.1 and is omitted.
@@
Fix $\alpha\in\mathbb{R}$. Then a given $\Omega\subset\mathbb{R}^{d}$ is diagonally determined with $\mathrm{diag}(\Omega)=[a,b]$ and vector $\mathbf{v}$} if and only if its translation $\Omega+\alpha\mathbf{1}$ is diagonally determined with diagonal $a+\alpha,b+\alpha]$ and vector $\mathbf{v}$}.
3. Least Chebyshev polynomial for a diagonally-determined set
In this section, and with $\Omega$ being a set in ${\mathbb{R}}^{d}$, we consider the problem of finding a least polynomial $P^{*}\in\Pi_{n}^{*}$ such that
@@ 3. Least Chebyshev polynomial for a diagonally-determined set
[…]
Then $P(t,\ldots,t)=t^{n}+q_{n-1}(t)$, where $q_{n-1}(t)=Q(t,\ldots,t)$ is a univariate polynomial of degree at most n1n-1 in the single variable $t$. Hence,
@@
where we have used a classical result in one variable (e.g. [12, Theorem 2.1]), and equality in the last inequality is attained by choosing $q_{n-1}$ such that $t^{n}+q_{n-1}(t)$ is the rescaled Chebyshev polynomial in the right-hand side of (3.2), when restricted to ${\mathbf{x}}=(t,...,t)$ ($t\in{\mathbb{R}}$).
@@
The minimal value in (3.1) does not depend explicitly on the dimension $d$, but the diagonal $[a,b]$ may. For example, if $\Omega$ is the unit Euclidean ball in ${\mathbb{R}}^{d}$, then $a,b]=[-1/\sqrt{d},1/\sqrt{d}]$.
@@
The simplex $\triangle^{d}=\{\,{\mathbf{x}}\in{\mathbb{R}}^{d}\;:\;0\leq x_{1}\leq x_{2}\leq\ldots\leq x_{d}\leq 1\,\}$ is diagonally-determined with ${\mathbf{v}}=(1,0,\ldots,0)$, and $\mathrm{diag}(\triangle^{d})=[0,1]$. Hence, by Theorem 3.1,
@@
Returning to Example 2.7, where $\Omega$ is the unit ball for the owl norm, we have [a,b]=[1/W,1/W][a,b]=[-1/W,1/W] and ${\mathbf{v}}={\mathbf{w}}/W$, so that, by Theorem 3.1, the least Chebyshev polynomial of degree $n$ becomes
@@
We point out that the requirement $\sum_{{\alpha}\in{\mathbb{N}}^{d}_{n}}a_{\alpha}=1$, is susceptible to an affine change of variables other than a translation. For instance, the triangle $\hat{\triangle}^{2}=\{\,{\mathbf{x}}:x_{1}\geq 0,x_{2}\geq 0,x_{1}+x_{2}\leq 1\,\}$, is a mirror image of $\triangle^{2}$ under $x_{2}\mapsto 1-x_{2}$. However, under this affine change of variables, the leading monomial $a_{0}x_{1}^{2}+a_{1}x_{1}x_{2}+a_{2}x_{2}^{2}$ of a polynomial in $\Pi_{2}^{*}$ becomes, $a_{0}x_{1}^{2}-a_{1}x_{1}x_{2}+a_{2}x_{2}^{2}$, which is, in general, no longer an element of $\Pi_{2}^{*}$.
@@
$\Omega_{1}$ $:=$ $\{(x,x)\;|\;x\in[-1,1]\}$
$\Omega_{2}$ $:=$ $\{(0,x)\;|\;x\in[-\sqrt{2},\sqrt{2}]\}$
$\Omega_{3}$ $:=$ $\{(x,-x)\;|\;x\in[-1,1]\}.$